# Local Thermal Operations and Classical Communication

### v1

@@ 3 Framework of Local Thermal Operations and Classical Communication
Having presented all the necessary foundational ideas and concepts, we may proceed to introduction of the central idea of this work, combining spatially separated parties with thermal restrictions on the operations that they can effect – Local Thermal Operations and Classical Communication (LTOCC).
@@ 3.1 Local thermal operations and preservation of product Gibbs state
Let us consider a scenario in which Alice and Bob have two separate physical subsystems and each one of them is able to perform thermal operations. Since two subsystems are physically separated, we assume that the interactions between them are negligible. Otherwise, the parties might leverage those interactions, effectively treating two subsystems as one joint system.
Therefore the total Hamiltonian is a sum of local Hamiltonians $H=H_{A}+H_{B}$. Hence the thermal state of the entire system is a product of thermal states on each subsystem ${\gamma}={\gamma}_{A}\otimes{\gamma}_{B}$. This assumption, natural for separated subsystems, results in a crucial property: The product of any local thermal operations is thermal as well.
@@ 3.2 Classical communication and Landauer cost of measurement
The final ingredient one requires before going into the framework of LTOCC is the problem of local measurements and classical communication between the parties, as there are certain subtleties that need to be taken into account, especially the energetic cost of measuring the system in the energy eigenbasis.
@@
The energy-incoherent quantum states $\rho=\sum_{i}p_{i}\;|E_{i}\rangle\langle E_{i}|$ can be interpreted as a probabilistic combination or a statistical ensemble of pure Hamintonian eigenstates. The measurement collapses it to a random eigenstate, thus one round of measurement is not thermal – it can project a Gibbs state onto a high energy eigenstate, resulting in an energetic gain. However, statistically, the measurements leave diagonal states invariant. Thus while performing measurement on consecutive states form a Gibbs ensemble, with respect to energy eigenbasis, $N$ times the disturbance effect of measurements vanishes on average as we go into the limit of very large number of rounds, $N\rightarrow\infty$.
The second element is the transfer and storage of information. The processing of information, according to the Landauer principle [55, 56], imposes a fundamental thermodynamic cost. Fortunately, this cost is associated with the erasure of information, which would happen between rounds or runs of protocol, on classical computers used for communication. Thus assuming that classical computers are well separated from the quantum system (except at the moment of measurement), one can disregard the contribution of Landauer cost. However, we may also limit ourselves to a finite sample and state, that after $N$ measurements the cost of erasing the memory can be estimated as $Nk_{B}T\log d$, where $d$ is a dimension of the quantum system, incurred on the side of the classical memory. After such erasure states from consecutive rounds are indistinguishable constituting the Gibbs ensemble. Thus statistically it is indeed preserved.
@@
Finally, because statistically neither thermal operations nor measurement in energy eigenbasis can create any coherences, description by probability distributions is complete, and instead of quantum channels, we can continue our discussion using stochastic maps.
@@ 3.3 Framework outline
First, let us state the minimal requirements for any action that Alice and Bob could perform while preserving respective local equilibria. Let $M_{ij,kl}$ be a stochastic map acting on two systems [^1] In what follows, unless stated otherwise, all stochastic tensors and matrices depend, implicitly, on the inverse temperature $\beta$.. Since neither Bob nor Alice can do anything to change the Gibbs state on either of the subsystems we have:
$\forall_{k}:\sum_{i}M_{ij,kl}=:B_{j,l}^{(1,k)};\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{l}:\sum_{j}M_{ij,kl}=:B_{i,k}^{(2,l)}\leavevmode\nobreak\ \leavevmode\nobreak$ (13)
where $B^{(1,k)}$, $B^{(2,l)}$ are thermal operations for all values of $k$ and $l$ respectively. In the above equation $k,l$ are input and $i,j$ are output indexes corresponding respectively to Alice’s and Bob’s subsystems. This can also be stated in terms of input and output states as
$\forall_{k}:\sum_{i,l}M_{ij,kl}\gamma_{l}=\gamma_{j};\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{l}:\sum_{j,k}M_{ij,kl}\gamma_{k}=\gamma_{i}\leavevmode\nobreak\ .$ (14)
We will call any channel, that satisfies those two properties locally Thermal Operation ($\ell$TO).
Now we proceed with the discussion of the actual scheme, concerned with what Alice and Bob can perform using local thermal operations, measurements and communication. After each of them receives their own part of the state, Alice can perform a measurement on her subsystem and send the information to Bob who, depending on the received information, can perform different thermal operations. If we stack Bob operations conveniently in one tensor $T_{jkl}$, the discussed example corresponds to the stochastic matrix of the form
$M_{ij,kl}=\delta_{i,k}T_{jkl}\leavevmode\nobreak\ ,$ (15)
since the ensemble of Alice’s states remains unchanged. In parallel to Alice, Bob could also perform a measurement of his input state, and send the information to Alice, who can also perform stochastic operations depending on the received information. Because measurement in the energy eigenbasis does not alter the ensembles of received states we can describe this joint operation as:
$M_{ij,kl}=T_{ikl}^{(1)}T_{jkl}^{(2)}$ (16)
which we will call one round (without feedback) Local Thermal Operations and Classical communication (LTOCC). The minimal requirements, applied on (16) results in
$\forall_{l}\sum_{k}T_{ikl}^{(1)}{\gamma_{E}}_{k}={\gamma_{E}}_{i}\;,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{k}\sum_{l}T_{jkl}^{(2)}{\gamma_{E}}_{l}={\gamma_{E}}_{j}\;.$ (17)
@@
This framework can be easily extended to multiple rounds without memory
$M=\prod_{a}M^{(a)}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ \leavevmode\nobreak\ M_{ij,kl}^{(a)}=T_{ikl}^{(1,a)}T_{jkl}^{(2,a)}$ (18)
or with memory, which would correspond to the growing number of indices for tensors in consecutive rounds, to describe dependencies on past measurement results. Thus replacing indices $k,l$ with multi-indices $\mathbf{k}=\quantity(k_{1},\ldots,k_{n}),\,\mathbf{l}=\quantity(l_{1},\ldots,l_{n})$ in the $n$-th round.
Another possible extension is the introduction of shared randomness (LTOCC+R). Alice and Bob can share a random variable $\lambda_{\alpha}$ and perform different LTOCC depending on its value:
$M_{ij,kl}=\sum_{\alpha}\lambda_{\alpha}T_{ikl}^{(1,\alpha)}T_{jkl}^{(2,\alpha)}$ (19)
Finally, we can impose an additional symmetry requirement. namely that if Alice received a thermal state as an input, then the result of Bob’s action would also be thermal and vice versa
$\forall_{k}\sum_{l}T_{ikl}^{(1)}{\gamma_{E}}_{l}={\gamma_{E}}_{i}\;,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{l}\sum_{k}T_{ikl}^{(1)}{\gamma_{E}}_{k}={\gamma_{E}}_{j}\;,$ (20)
and analogously for $T_{jkl}^{(2)}$.
We call any stochastic tensor satisfying both properties (20) a bi-thermal tensor. The above scheme constructed from a bi-thermal tensor gives one round of Symmetric Thermal Operations and Classical Communication ($\mathcal{S}$LTOCC).
In this work, we focus on the building block of this framework, the (bi-)thermal tensor $A$, discussing its properties, the structure of the set of such tensors and the possibility of obtaining outcomes for given input distributions.
@@
Figure 3: Diagram of inclusions of all above-discussed classes of quantum protocols. Each frame corresponds to a different number of rounds in the protocol.
4 Detailed results on $\mathcal{S}$LTOCC
In this section, we shift our attention to the building block of the presented framework – thermal tensor. Particularly we focus on bi-thermal tensors, with the additional requirement of symmetry, to present its wide range of properties even in such restricted case. We start the discussion by considering the geometry of bi-thermal tensors’ set to unveil the extremal operations.

### v2

@@ 4 Framework of Local Thermal Operations and Classical Communication
Having presented all the necessary foundational ideas and concepts, we may proceed to the introduction of the central idea of this work, combining spatially separated parties with thermal restrictions on the local operations that they can effect, with classical communication between them – Local Thermal Operations and Classical Communication (LTOCC).
@@ 4.1 Local thermal operations and preservation of product Gibbs state
Let us start by considering a scenario in which Alice and Bob have two separate physical subsystems and each of them is able to perform thermal operations locally. Since two subsystems are physically separated, we assume that the interactions between them are negligible. Otherwise, the parties might leverage those interactions, effectively treating two subsystems as one joint system.
Therefore, the total Hamiltonian can be expressed as a sum of local Hamiltonians $H=H^{(A)}+H^{(B)}$. Hence, the thermal state of the entire system is a product of thermal states on each subsystem, ${\gamma}^{(AB)}={\gamma}^{(A)}\otimes{\gamma}^{(B)}$ together with their corresponding local inverse temperatures $\beta^{(A)}$ and $\beta^{(B)}$. This assumption, natural for separated subsystems, results in a crucial property: the product of any local thermal operations is (semilocal) thermal as well.
@@ 4.2 Classical communication and Landauer cost of measurement
The final ingredient one requires before going into the framework of LTOCC is the problem of local measurements and classical communication between the parties, as there are certain subtleties that need to be taken into account, especially the thermodynamic cost of measuring the system in the energy eigenbasis. Such measurements can be emulated using the SLTO framework given memory erasure after each round of measurement and forwarding of information, which we demonstrate explicitly in D. This result is similar in spirit to the incoherently conditioned thermal operations as a subset of TO, as demonstrated in [58]. Below, we lay down more heuristic arguments and point to a possible hidden cost connected with Landauer erasure.
@@
The energy-incoherent quantum states $\rho=\sum_{i}p_{i}\;\outerproduct{E_{i}}{E_{i}}$ | can be interpreted as a probabilistic combination or a statistical ensemble of pure Hamiltonian eigenstates. The projective measurement onto the energy eigenbasis collapses it to a random eigenstate |$\ket{E_{i}}$ with probability $𝐸𝑖p_{i}=\expectationvalue{\rho}{E_{i}}$, thus single-shot measurement is not a thermal operation – it can project a Gibbs state onto a high or low energy eigenstate, resulting in a gain of free energy, otherwise inaccessible by thermal operations alone. However, statistically, incoherent measurements leave diagonal states invariant $\sum_{i}|E_{i}\rangle\langle E_{i}|\langle E_{i}|\rho|E_{i}\rangle=\sum_{i}p_{i}\;\outerproduct{E_{i}}{E_{i}}=\rho$. Thus while performing measurement on consecutive states from a Gibbs ensemble, with respect to energy eigenbasis $N$ times, the discrepancy between an ensemble of measurement outcomes and initial distribution vanishes on average as we go into the limit of a large number of rounds, $N\rightarrow\infty$.
The second element is the transfer and storage of information. According to Landauer’s principle [59, 60], erasing information in a classical memory incurs a fundamental thermodynamic cost. In our analysis, we assume that the classical computers used to store measurement outcomes are thermodynamically isolated from the quantum system, except for the moment of saving the outcome. Thus, although the erasure of classical memory (required between rounds or runs of the protocol) carries a cost of approximately $Nk_{B}T\log d$ for $N$ measurements (with $d$ being the dimension of the quantum system), this cost does not affect the quantum system itself. Here, “erasure” refers solely to the resetting of the classical memory used in the LOCC protocol, not to any direct manipulation of the quantum state. Consequently, the Landauer cost can be disregarded when considering the quantum system’s thermodynamics.
Additionally, we note that arguments made by Watanabe and Takagi in [58] do not translate to our setting, as Alice and Bob do not have access to shared bath, and both systems may be in different temperatures. Thus, in order to formalise heuristic considerations put forward above we explicitly consider measurement in energy eigenbasis as a channel $\Phi:\mathcal{H}_{A}\mapsto\mathcal{H}_{A}\otimes\mathcal{M}$ represented by a set of operators $K_{i}=\quantity(\ket{E_{i}}\otimes\ket{i}_{\mathcal{M}})\bra{E_{i}}$ | for all $i$, which imprint the detected information about the energy level $E_{i}$ on an auxiliary (classical) memory system $\mathcal{M}$ and assume that energies are non-degenerate. It is straightforward to see that after discarding memory the map returns pinched state $𝐸𝑖\text{Tr}_{\mathcal{M}}\quantity[\Phi(\rho)]=\mathcal{P}(\rho)=\sum_{i}\expectationvalue{\rho}{E_{i}}\outerproduct{E_{i}}{E_{i}}$ |. Pinching map, in general, preserves all energy-incoherent states and, in particular, the Gibbs state. Furthermore, the use of classical information obtained from measurement to apply a classically-conditioned map is described by channel $\Xi:\mathcal{M}\otimes\mathcal{H}_{B}\mapsto\mathcal{M}\otimes\mathcal{H}_{B}$ defined as $\Xi(\mu\otimes\sigma)=\sum_{i}\langle i|\mu|i\rangle|i\rangle\langle i|\otimes\Xi_{i}(\sigma)$, with $\quantity{\Xi_{i}:\mathcal{H}_{B}\mapsto\mathcal{H}_{B}}$ being a collection of CPTP maps. From this discussion one can infer that any further manipulation of subsystems $\mathcal{M}$ and $\mathcal{H}_{B}$ cannot alter the marginal belonging to the system $\mathcal{H}_{A}$:
$\text{Tr}_{B,\mathcal{M}}\quantity[(\mathbb{I}_{A}\otimes\Xi)\,\circ\,(\Phi\otimes\mathbb{I}_{B})\quantity(\rho\otimes\sigma)]=\text{Tr}_{\mathcal{M}}\quantity[\Phi(\rho)]=\mathcal{P}(\rho)=\rho.$
We highlight that the above consideration assumes free access to, updating and erasure of a memory register – which is far from obvious, but explained and justified in the prior paragraphs.
Furthermore, let us highlight that one could extend the possible measurements to more general POVMs, e.g. with operators of the form $K_{i}=\sum_{j}a_{ij}\outerproduct{E_{j}}{E_{j}}$ | such that for all $j$ $\sum_{i}a_{ij}^{2}=1$, while satisfying the Gibbs preservation condition. There are three points to be made in this context. Firstly, when restricted to energy-incoherent states, such operations can be easily achieved equivalently by classical post-processing on the memory side. The second, more important aspect is that such measurements would in general provide partial coherence and/or entanglement preservation while giving only imperfect information about the system. Thirdly, althou such energy incoherent measurements preserve Gibbs state, in case of general quantum states it may not be strong enought restriction to consider it “thermal”, due to the distinction between thermal operations and Gibbs-preserving operations, which emerge on the quantum. However, due to their connection to coherence such measurements fall outside of the scope of this manuscript, and will be considered in detail in a future work [61].
@@
Finally, because statistically neither (conditional) thermal operations nor measurement in energy eigenbasis can create any coherences, the description of energy incoherent states by probability distributions is complete, and instead of quantum channels, we can continue our discussion using stochastic maps.
@@ 4.3 Framework outline
In what follows for simplicity of exposition, we will focus on the bipartite scenario, with extension of the general multipartite definitions following in a similar manner.
We start our discussion of local thermal operations aided by classical communication with one round scenario restricted by operational consideration, where Alice and Bob are restricted to thermal operations in each of their respective subsystems, but can also perform incoherent measurements. We will denote local Gibbs state as $\gamma^{(A)}$ and $\gamma^{(B)}$, respectively.
The protocol proceeds as follows: in the first step, Alice can perform a measurement on her subsystem and send the information to Bob who, depending on the received information, can perform different thermal operations; in other words, Bob performs local thermal operations conditioned on the information obtained by Alice from measurement. If we stack Bob operations conveniently in one stochastic tensor $T_{jkl}^{(B)}$, the discussed scheme corresponds to a stochastic matrix of the form
$M_{ij,kl}=\delta_{ik}T_{jkl}^{(B)}\leavevmode\nobreak\ ,$ (15)
since the ensemble of Alice’s states remains unchanged; keep in mind that the superscript (B) refers to Bob, the party performing thermal operations. Combining that with possible local thermal operation $\Lambda^{(A)}$ on Alice’s side
$M_{ij,kl}=\Lambda_{ik}^{(A)}T_{jkl}^{(B)}$ (16)
[…]
$M_{ij,kl}=T_{ikl}^{(A)}\Lambda_{jl}^{(B)}\leavevmode\nobreak$ (17)
From the thermality of each conditioned thermal operation in (16) and (17) one obtains thermal restrictions for the tensors $T$:
$\forall_{l}\sum_{k}T_{ikl}^{(A)}\gamma_{k}^{(A)}=\gamma_{i}^{(A)}\;,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{k}\sum_{l}T_{jkl}^{(B)}\gamma_{l}^{(B)}=\gamma_{j}^{(B)}\;.$ (18)
@@
The framework of local thermal operations and classical communication can be easily extended to multiple rounds without memory
$M=\prod_{a}M^{(\alpha)}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ \leavevmode\nobreak\ M_{ij,kl}^{(\alpha)}=\Lambda_{ik}^{(A,\alpha)}T_{jkl}^{(B,\alpha)}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ T_{ikl}^{(A,\alpha)}\Lambda_{jl}^{(B,\alpha)}\leavevmode\nobreak\ ,$ (19)
with $\alpha$ indexing the consecutive rounds. We denote all operations realizable using $n$-round protocol of the form (19) as $\text{LTOCC}_{n}$. A schematic representation of this multiround protocol is provided in Fig. 3(a).
Further extension can be made by allowing both Alice and Bob to retain a memory of measurement results from previous rounds, referred to as LTOCC+M.
$\mathbf{M}\quantity(\mathbf{p}\otimes\mathbf{q})=\text{Tr}_{\mathcal{M}_{1},\ldots,\mathcal{M}_{n}}\quantity(\quantity[\bigcirc_{\alpha}\mathbf{M}^{(\alpha)}]\quantity(\mathbf{p}\otimes\mathbf{q})).$ (20)
Each round consists of measurement on one subsystem combined with local thermal operations conditioned by all previous measurements, ie:
$r_{ij,m_{1}\ldots m_{\alpha}}^{\alpha}=\left[\mathbf{M^{(\alpha)}}(\mathbf{r}^{(\alpha-1)})\right]_{ijm_{1}\ldots m_{\alpha}}=\sum_{kl}{T^{(A,\alpha)}}_{i}^{k;\,m_{1}\ldots m_{\alpha-1}}\delta_{k,m_{\alpha}}\;{T^{(B,\alpha)}}_{j}^{l;\,m_{1}\ldots m_{\alpha-1}m_{\alpha}}\;r_{kl,m_{1}\ldots m_{\alpha-1}}^{(\alpha-1)}$ (21)
where $m_{i},\ldots,m_{\alpha}$ are indices related to the memory registers, $k,l$ are system-related input state indices, and $i,j$ are system-related output-state indices. Here, $\delta_{k,m_{\alpha}}$ is a measurement on Alice’s subsystem, recorded on a new classical register $m_{\alpha}$. This is followed by thermal operation ${T^{(A)}}_{i}^{k;\,m_{1}\ldots m_{\alpha-1}}$ conditioned on all previous measurement results $m_{1}\ldots,m_{\alpha-1}$. Furthermore ${T^{(B)}}_{j}^{l;\,m_{1}\ldots m_{\alpha-1}m_{\alpha}}$ is a thermal operation on Bob’s side conditioned by the measurement’s outcome in previous rounds $m_{1},\ldots,m_{\alpha-1}$ and new measurement result on the Alice side $m_{\alpha}$. We provide a diagrammatic representation of this protocol in Fig.  3(b). Additional detailed discussion of 1- and 2-round LTOCC protocols with and without protocols is provided in C.
(a) $\text{LTOCC}_{2}$
@@
@C=1.0em @R=1.0em @!R
[…]
(b) $\text{LTOCC}_{2}+\text{M}$
[…]
@C=1.0em @R=1.0em @!R
[…]
Figure 3: Two-round LTOCC with and without memory: In both versions of $\text{LTOCC}_{2}$ protocol first round consists of Alice measuring her subsystem, storing the result in the memory register $m_{1}$, which is then sent to Bob to implement conditional thermal operation, represented by the thermal tensor $T^{(B,1)}$, which is followed with Alice implementing a local post-processing TO $\Lambda^{(A,1)}$. In the no-memory case (a) in the second round the memory register $m_{1}$ is overwritten with Bob’s measurement of his system and the operations are continued with the roles swapped. For memory case (b) the register $m_{1}$ is retained and Bob stores his result in the new register $m_{2}$. Subsequently, Alice performs operations conditioned on both $m_{1}$ and $m_{2}$, whereas Bob post-processes his subsystem conditioned on $m_{1}$ only.
We supplement the exhibition of multiround LTOCC protocols with and without memory with two simple observations regarding inclusions.
Observation 1.
$\text{LTOCC}_{n}$(+M) forms a superset of $\text{LTOCC}_{n^{\prime}}$(+M) for any $n^{\prime}\leq n$.
Proof.
The observation follows by considering $\text{LTOCC}_{n+1}$(+M), where Bob measures in $n+1$-st round, and setting $T^{(A,n+1)}$ and $\Lambda^{(B,n+1)}$ ($T^{(B,n+1)}$) to unconditional identity operations, thus realising arbitrary $n$-round protocol. The claim follows by induction. ∎
Observation 2.
$\text{LTOCC}_{n}+\text{M}$ forms a superset of $\text{LTOCC}_{n}$.
Proof.
The observation follows by setting $T^{(A,n)}$ to be conditioned nontrivially only on $m_{n}$ and $T^{(B,n)}$ to be unconditioned in rounds when Bob measures, and similarly for the rounds when Alice measures. ∎
Note that we do not provide the arguments for proper inclusion relations between protocols with and without memory, as well as for protocols of different depths. Such a problem is nontrivial even for standard LOCC [46].
Utilising two-round LTOCC+M one can construct a particularly interesting example, in which Alice and Bob measure their respective subsystems in parallel, and then operate on their respective systems based on information received from their counterpart. It can be realised in two rounds as follows. The first round consists of Alice measuring her system and storing the result in memory register $m_{1}$ without performing any postprocessing $T^{(A,1)}:=\Lambda^{(A,1)}=\mathbb{I}$. Bob also does not perform any action $T^{(B,1)}=\mathbb{I}$ (unconditionally). When the second round begins, Bob measures his system and stores the result in a second memory register $m_{2}$. Then we set Alice operation $T^{(A,2)}$ to be conditioned on $m_{2}$ only and, similarly, Bob’s postprocessing $T^{(B,2)}$ on previously stored $m_{1}$. Evaluating (21) according to the above description leads to a compact 1-step transition matrix
$M_{ij,kl}=T_{ikl}^{(A)}T_{jkl}^{(B)}.$ (22)
In a slight abuse of terminology, we refer to this particular restricted scenario as 1-round parallel LTOCC, as it does not involve any feedback. This procedure is depicted in Fig.  4.
[…]
@C=1.0em @R=1.0em @!R
[…]
Figure 4: Parallel LTOCC as a special case of $\text{LTOCC}_{2}+\text{M}$: Parallel protocol can be implemented as a variant of 2-round protocol with memory by limiting first round of the protocol to measurement only and then conditioning the local operations on the memory register of the other party – Alice on Bob’s measurement and Bob on Alice’s measurement.
Finally, one may further extend the framework by introducing shared randomness (LTOCC+R). Alice and Bob can share a random variable $\lambda_{\alpha}$ and perform different LTOCC depending on its value:
$M_{ij,kl}=\sum_{\alpha}\lambda_{\alpha}\Lambda_{ik}^{(A,\alpha)}T_{jkl}^{(B,\alpha)}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ M_{ij,kl}=\sum_{\alpha}\lambda_{\alpha}T_{ikl}^{(A,\alpha)}\Lambda_{jl}^{(B,\alpha)}$ (23)
Below, we demonstrate that shared randomness provides an actual advantage over LTOCC without shared randomness, showing that there exist operations which can be realised by LTOCC+R and not by LTOCC.
Theorem 6.
Local Thermal Operations and Classical Communication (LTOCC) form a strict subset of LTOCC with randomness (LTOCC+R).
Proof.
To show that LTOCC is a proper subset of LTOCC+R we first show that LTOCC is not a convex set and then that LTOCC+R is its convex hull. Hereafter, for the sake of clarity, we focus on one round of the protocol, however, the presented arguments hold for any number of rounds.
Let us first note that, similarly to thermal operation $\Lambda$, a set of thermal tensors $T_{ikl}$ forms a convex polytope since it is defined by linear relations together with the linear inequalities $0\leq T_{ikl}\leq 1$. Thus each tensor $T_{ikl}$ can be decomposed as a convex combination of extremal tensors $T_{ikl}=\sum_{\alpha_{1}}t_{\alpha_{1}}T_{ikl}^{\alpha_{1}}$ with some weights $t_{\alpha_{1}}$; similarly for $\Lambda_{ik}=\sum_{\alpha_{2}}u_{\alpha_{2}}\Lambda_{ik}^{\alpha_{2}}$. However, the set of LTOCC maps $M_{ij,kl}=\Lambda_{ik}^{(A)}T_{jkl}^{(B)}$ is not convex, since each map can be decomposed into extremal LTOCC using product distribution $M_{ij}=\sum_{\alpha_{1},\alpha_{2}}t_{\alpha_{1}}u_{\alpha_{2}}\Lambda_{ik}^{\alpha_{2}}T_{jkl}^{\alpha_{1}}=:\sum_{\alpha_{1},\alpha_{2}}t_{\alpha_{1}}u_{\alpha_{2}}M_{ij,kl}^{\alpha_{1},\alpha_{2}}$. Since the product distributions form proper subsets of all two-dimensional distributions we infer some combinations of extremal maps are not in $LTOCC$, thus it is not a convex set.
[…]
Similar statements follow mutatis mundis when considering multi-round protocols with the same number of rounds $n$ both with and without memory, when comparing variants with and without randomness (R).
Finally, we present the main property of most of the introduced protocols, which is simultaneously a main restriction of their capabilities:
Theorem 7.
Local thermal Operations and Classical communication with randomness (LTOCCn +R) (with projective energy-incoherent measurements), form a subset of semilocal thermal operations (SLTO) (even for non energy-incoherent states).
Proof.
Proof of this theorem is presented in D. ∎
We stress that the following theorem describes only protocols without memory. If we allow the use of outcomes of some measurements in multiple rounds, the class of obtained protocols might be noticeably more powerful, as presented in the C.
Corollary 3 (Thermomajorisation for LTOCC).
Consider two systems $A$ and $B$ with local Gibbs states $\gamma^{(A)}$ and $\gamma^{(B)}$, respectively. Given an input state $\mathbf{p}_{AB}$ and output $\mathbf{q}_{AB}$, the transition can be achieved using LTOCCn +R only if
$\mathbf{p}\underset{\gamma_{AB}}{\succ}\mathbf{q}$ (24)
where the thermomajorization is taken with respect to $\gamma_{AB}=\gamma^{(A)}\otimes\gamma^{(B)}$. In particular, the Gibbs state is preserved
$M(\gamma^{(A)}\otimes\gamma^{(B)})=\gamma^{(A)}\otimes\gamma^{(B)}$
We stress that the above corollary describes the implication in only one direction. The implication in the opposite direction would imply that LTOCC $=$ SLTO.
Before finishing this section, we can impose an additional symmetry requirement. Consider a scenario in which Alice’s and Bob’s thermal states coincide $\gamma^{(A)}=\gamma^{(B)}=:\gamma$, and a parallel LTOCC protocol in which if Alice received a thermal state as an input, then the result of Bob’s action would also be thermal and vice versa
$\forall_{k}\sum_{l}T_{ikl}^{(A)}\gamma_{l}=\gamma_{i}\;,\leavevmode\nobreak\ \leavevmode\nobreak\ \forall_{l}\sum_{k}T_{ikl}^{(A)}\gamma_{k}=\gamma_{j}$ (25)
and analogously for $T_{jkl}^{(B)}$. Thus Alice and Bob could exchange their tensor without going outside of the framework of LTOCC. We call any stochastic tensor satisfying both properties (25) a bithermal tensor. The above scheme constructed from a bithermal tensor gives one-round Symmetric Thermal Operations and Classical Communication ($\mathcal{S}$LTOCC).
Theorem 8.
Symmetric Local Thermal Operations and Classical Communication with $n$ rounds ($\mathcal{S}$LTOCCn) form a strict subset of $2n$-rounds Local Thermal Operations and Classical Communication (LTOCC2n).
Proof.
The inclusion follows from the fact that each round $\mathcal{S}$LTOCC, by definition, can be implemented in two rounds of LTOCC. The simplest proof of the proper inclusion is based on the observation that both $\mathcal{S}$LTOCC output marginals need to be Gibbs wherever the Gibbs state appears as one of the inputs. On the other hand, this is not the case for all LTOCCs, which can be presented using the identity protocol in which all (conditional) terminal operations are just identity. ∎
An alternative and more insightful argumentation requires the demonstration that $\mathcal{S}$LTOCC’s building block – bithermal tensor forms a proper subset of the set of thermal tensors. Here we focus once again on one round of $\mathcal{S}$LTOCC for convenience notation, however, the presented argument is general. To do so, we note that $T_{ijk}$ in $\mathcal{S}$LTOCC is restricted by more linear conditions than in LTOCC (25). Furthermore, each equation defining thermal tensors $\forall_{l}\sum_{k}T_{ikl}^{(A)}\gamma_{k}=\gamma_{i}$ does not include the elements from different layers of the tensor (different values of $l$), whereas the second set of equalities does so. Hence, one set of equations cannot impose the other, so the additional equalities provide further restrictions. The proper inclusion of the bithermal tensors’ set inside the thermal tensors’ set for local dimension $d=2$ is also presented in the next section.
Theorem 9.
Symmetric Thermal Operations and Classical Communication ($\mathcal{S}$LTOCC) form a strict subset of $\mathcal{S}$LTOCC with randomness.
Proof.
The proof follows mutatis mutandis the line of reasoning of Theorem 6. ∎
Summary of inclusion relations between variants of LTOCC model with different amounts of memory and control, as discussed in this section, is presented pictorially in Fig. 5.
Finally, we put forward a conjecture concerning the relation between LTOCC and SLTO in general.
Conjecture 1.
Local thermal operations and classical operations with an arbitrary number of rounds and shared randomness yield a set of operations, which closure is equivalent to the semilocal thermal operations when we restrict both classes to energy-incoherent states,
$\overline{\bigcup_{n=1}^{\infty}(\text{LTOCC}_{n}+\text{R})}=\text{SLTO}.$ (26)
Should the above conjecture, or its stronger version without shared randomness or extension to genuinely quantum states, hold true, it would have interesting consequences, especially taking into account that considerations in [33] have been carried out appealing only energy-incoherent states – finite-size engine operating with Carnot efficiency could be realised with a (potentially finite-round) LTOCC protocol without appealing to its quantum properties. Furthermore, in LTOCCn +R all interactions between subsystems and the extension between thermal baths are purely classical, which is not the case for SLTO. Hence, the above result would indicate that at least for energy-incoherent states, the introduction of coherent interactions does not increase the set of possible operations.
[…]
Figure 5: LTOCC inclusion hierarchy: Each frame corresponds to a different number of rounds in the protocol. In general, variants with shared randomness (+R) contain the ones not including it (Theorems 6 and 9), $n$-round protocols contain all $n^{\prime}$-round protocols for $n^{\prime}\leq n$ (Observation 1), protocols with memory (+M) contain the ones without it (Observation 2), and protocols without memory (including all one round protocols) form a subset of SLTO (Theorem 7).
5 Basic building blocks – thermal and bithermal tensors
In this section, we shift our attention to the new building block of the presented LTOCC framework – thermal tensor – since the properties of thermal operations have already been widely studied, e.g. [62]. We start with the observation that the layers of thermal tensors are independent Gibbs-preserving stochastic operations. Therefore the space of possible thermal tensors as a whole can be summarised easily as $d$-th tensor power of the space of Gibbs-preserving matrices, which has already been studied in detail [55, 62, 45]. Thus, solutions to certain problems, like extremal operations, are straightforward, given that they are known for Gibbs-preserving matrices.
Taking this into account, we narrow down a majority of the discussion to a mathematically interesting and unexplored subset constituting the building blocks of $\mathcal{S}$LTOCC – bithermal tensors, with the additional requirement of symmetry, to present its wide range of properties in such a restricted case. Additionally, we note that this set is interesting in its own right due to its semblance to already explored problems – bithermal and tristochastic tensors bear a similar relation as thermal-stochastic and bistochastic matrices. We start the discussion by considering the geometry of the bithermal tensors’ set particularly to unveil the extremal operations – a thermal equivalent of study of Birkhoff-like polytope for tristochastic tensors.